You are creating identical coupon booklets to sell for a fundraiser. You have 32 drry-cleaning coupons, 80 coupons for a local restaurant, and 47 movie theater coupons. Can you create multiple, identical coupon booklets using all of the coupons? You find one additional movie theater coupon. What is the greatest number of identical coupon booklets you can make?
Answer by vleith(2983)

47 is a prime number. Since the other two value are not multiples of 47, the original numbers can't be divvied into multiple books
If you now have 48 movies tickets, you have a chance.
factor each value given into it prime factors
32 = 2* 2 * 2 * 2 *2
80 = 2* 2 * 2 * 2 * 5
48 = 2* 3 * 2 * 2 * 2
each of these numbers has the factor 2 in it 4 times = 16
you can make 16 booklets. Each booklet has 2 dry cleans, 5 restaurant, and 3 movie coupons.
Get it?